# Effects of particle-hole fluctuations on the superfluid transition in two-dimensional atomic Fermi gases

### v1

@@ Abstract
Proper treatment of the many-body interactions is of paramount importance in our understanding of strongly correlated systems. Here we investigate the effects of particle-hole fluctuations on the Berezinskii-Kosterlitz-Thouless (BKT) transition in two-dimensional Fermi gases throughout the entire BCS-BEC crossover. We include self-consistently in the self energy treatment the entire particle-hole $T$ matrix, which constitutes a renormalization of the bare interaction that appears in the particle-particle scattering $T$ matrix, leading to a screening of the pairing interaction and hence a dramatic reduction of the pairing gap and the transition temperature. The BKT transition temperature $T_{\text{BKT}}$ is determined by the critical phase space density, for which the pair density and pair mass are determined using a pairing fluctuation theory, which accommodates self-consistently the important self-energy feedback in the treatment of finite-momentum pairing fluctuations. The screening strength varies continuously from its maximum in the BCS limit to essentially zero in BEC limit. In the unitary regime, it leads to an interaction-dependent shift of $T_{\text{BKT}}$ towards the BEC regime. This shift is crucial in an attempt to explain experimental data quantitatively, which often depends on the interaction strength. Our findings are consistent with available experimental results in the unitary and BEC regimes and with quantum Monte Carlo simulations in the BCS and unitary regimes.
@@ I Introduction
Strongly correlated systems constitute the main challenge and are thus at the heart and frontier of in condensed matter physics. Due to the low dimensionality, fluctuations in two dimensions (2D) are usually strong, positioning the system in the strongly correlated regime, for which a proper treatment of interaction effects is of paramount importance, in order to understand various experiments and physical phenomena at a quantitative level.
@@
In the absence of long range order a la the Mermin-Wagner theorem [1], phase transition in 2D is in general of the Berezinskii [2], Kosterlitz and Thouless (BKT) [3] type. Indeed, superfluid transitions in ultracold atomic Bose gases have been experimentally found to exhibit a strong BKT nature in 2D [4, 5, 6]. BKT transitions have been of wide interest since the most important high $T_{c}$ superconductors are quasi-2D layered materials [7], for which the low dimensionality is favorable for the formation of the wide-spread pseudogap phenomena, which are arguably attributable to strong fluctuations [8, 9]. It is also the model commonly used for describing the superconducting transition in thin films [10, 11]. Recent exciting discoveries of (quasi-)2D superconductors [12, 13, 14, 15] have added to the interest in 2D phase transitions. The BKT nature of the superfluid transition in 2D atomic Fermi gases has also been experimentally confirmed [16, 17]. There are also, albeit not many, theoretical studies of the BKT superfluid transition in a 2D Fermi system for both weak and strong pairing interactions, e.g., Ref. [18, 19].
@@
In this paper, we report our study on the important yet often neglected effect of the particle-hole fluctuations on 2D Fermi gas superfluidity, to which similar studies has not been reported in the literature [19]. Similar to its 3D counterpart, we find that particle-hole fluctuations lead to a screening to the pairing interaction, causing a shift of the superfluid transition curve toward the BEC regime as a function of the pairing strength. This substantially suppresses $T_{c}$ and the gap in the BCS and unitary regimes, and has an important physical significance when comparing between theory and experiment.
@@
It is important to note that there has been a historical controversy surrounding 2D fermionic superfluids, dating back to Kosterlitz and Thouless [3], which primarily concerns observable signatures and the applicability of BKT physics. It is not straightforward to apply the BKT theory based on the XY model to fermionic superfluids. In the XY model, in which the superfluid density $n_{s}/m$ provides the phase stiffness, it is the phase fluctuations of the vortex type that dominate the destruction of the quasi-long range superfluid order and drive the system into a disordered normal state. Unlike the superfluids of true bosons, however, the superfluid density is also suppressed by the pair-breaking Bogoliubov quasiparticle excitations as the temperature $T$ increases. In addition, both the density and the mass of the bosons (i.e., fermion pair) now depend on the temperature and the pairing interaction strength, except in the deep BEC regime. Therefore, computing the superfluid density using a mean-field approximation would yield a constant value $n_{s}/m=n/m$ at $T=0$ (or $n/4m$ from the boson point of view, with $n_{B}=n/2$ and $m_{B}=2m$), independent of the interaction strength. (Following standard notations, here $n$, $m$, $n_{B}$, $m_{B}$ and $n_{s}$ denote total fermion number density, fermion mass, Cooper pair number density, Cooper pair mass, and superfluid number density, respectively.) At finite $T$, the pair-breaking effect due to quasiparticle excitations can be partly accounted for via solving the mean-field BCS gap equation, leading to a mean-field result of $(n_{s}/m)^{\text{BCS}}$. However, not being able to properly take care of the effective pair mass and pair number density is thus expected to give rise to an overestimate of the superfluid transition temperature $T_{\text{BKT}}$. Indeed, the fact $T_{\text{BKT}}$, as measured experimentally in 2D atomic Fermi gases, varies with interaction strength, reflects that both $n_{\text{B}}$ and $M_{\text{B}}$ vary with the interaction.
@@
The pair number density at given temperature and interaction is normally governed by the pair dispersion. The latter, or equivalently pair mass, can, in principle, be extracted from the pair propagator. However, there are not many calculations of the fermionic $T_{\text{BKT}}$ in the literature [20, 21, 22, 23], partly due to the inapplicability of the simple XY model directly to the Fermi system. The fact that $(n_{s}/m)^{\text{BCS}}$ decreases with temperature below $T_{\text{BKT}}$ in the BCS regime manifests the fact that amplitude fluctuations play an important role. Therefore, the superfluid transition in a 2D Fermi gas has a much richer physics than the simple XY model can capture. To account for the finite temperature and interaction effects, Wu et al [24] and Wang et al Ref. [18] determined $T_{\text{BKT}}$ using the critical phase space density criterion, based on a quantum Monte Carlo simulation [25] along with experimental support [17]. A pairing fluctuation theory [26] that was developed to address the pseudogap phenomena in 3D superconductors was used to determine the effective pair number density $n_{\text{B}}$ and pair mass $M_{\text{B}}$, which reflect the effects of both amplitude and phase fluctuations, governed by temperature and interaction strength. Nevertheless, this earlier work considered only the particle-particle channel of the $T$-matrix.
@@
In 2D, there have been no similar studies in the literature, however. It is the purpose of the present work to investigate the effect of particle-hole fluctuations on the atomic Fermi gas superfluidity in 2D. The low dimensionality further enhances the strong pairing fluctuations, which are already present when the interaction is strong. The very BKT nature of the transition requires that there is already a sizable pairing amplitude, i.e., (pseudo)gap, at $T_{\text{BKT}}$. Therefore, this necessitates the self-consistent inclusion of the self-energy feedback in (both the particle-particle and) the particle-hole $T$-matrices. Here we incorporate the particle-hole channel contributions to the pairing fluctuation theory [26, 33, 32], and thus go beyond previous studies [24, 18]. This pairing fluctuation theory includes self-consistently the finite-momentum pairing fluctuations in the single-fermion self energy, and has successfully addressed multiple high $T_{c}$ [26, 33, 34, 35] and atomic Fermi gas experiments [36, 37, 38, 39]. In particular, it features naturally a pseudogap in 3D unitary Fermi gases, which has been unequivocally corroborated by a recent experiment [40]. Following the previous work in 3D [32], we show that the particle-hole $T$-matrix serves as a renormalized pairing interaction, which is to appear in the $T$-matrix in the usual particle-particle channel. We find that the particle-hole fluctuations lead to a screening to the pairing interaction, and the inverse pairing interaction is effectively shifted by the temperature-dependent, angular-averaged (at two different levels) particle-hole susceptibility $\langle\chi_{\text{ph}}\rangle$, which approaches a negative constant, $m/2\pi$, in the BCS limit and increases gradually in the crossover regime toward zero in the BEC limit. In result, the particle-hole channel shifts the $T_{\text{BKT}}$ curve towards the BEC regime as a function of pairing strength. Furthermore, comparison shows that the inclusion of particle-hole channel leads to a better overall agreement between our calculated $T_{\text{BKT}}$ with the results from experiment and quantum Monte Carlo simulations.
@@ II.1 Overview of the Pair Fluctuation Theory without the Particle-Hole Channel
We consider a 2D Fermi gas with a short-range $s$-wave attractive interaction $V_{\mathbf{k},\mathbf{k}^{\prime}}=U<0$, as described by a generic grand canonical Hamiltonian [33], which includes pairing with a finite momentum $\mathbf{q}$. The free fermion dispersion is given by $\xi_{\mathbf{k}}=\epsilon_{\mathbf{k}}-\mu\equiv\mathbf{k}^{2}/2m-\mu$, with chemical potential $\mu$. The Fermi momentum is $\hbar k_{\text{F}}=\sqrt{2\pi n}\hbar$, and Fermi energy $E_{\text{F}}\equiv k_{\text{B}}T_{\text{F}}=\hbar^{2}k_{\text{F}}^{2}/2m$. As usual, we shall use the natural units, and set $\hbar=k_{\text{B}}=1$ and the volume to unity [26].
The fermions acquire self energy via pair binding and unbinding. Our approximated equations, derived via an equations of motion approach [41, 42], can be cast into a $T$ matrix formalism, with a mix of bare Green’s function $G_{0}(K)=(\mathrm{i}\omega_{l}-\xi_{\mathbf{k}})^{-1}$ and full Green’s function $G(K)$ in the pair susceptibility $\chi(Q)=\sum_{K}G(K)G_{0}(Q-K)$. Here we use the four-momentum notation, $K\equiv(\mathrm{i}\omega_{l},\mathbf{k})$ and $Q\equiv(\mathrm{i}\Omega_{n},\mathbf{q})$, $\sum_{K}\equiv T\sum_{l,\mathbf{k}}$ and $\sum_{Q}\equiv T\sum_{n,\mathbf{q}}$, with $\omega_{l}$ and $\Omega_{n}$ being Matsubara frequencies for fermions and bosons, respectively [43]. The $T$-matrix $t(Q)$ is given by $t(Q)=U/[1+U\chi(Q)]$, with the self energy
@@
The Thouless criterion (for superfluid transition in 3D) requires $t^{-1}(Q=0)=U^{-1}+\chi(0)=0$. In order to accommodate the absence of long range order in 2D, we generalize the Thouless criterion to allow for a very small but finite pair chemical potential $\mu_{\text{p}}$, $U^{-1}+\chi(0)=t^{-1}(0)\propto\mu_{\text{p}}$. In this way, $t(Q)$ is highly peaked around $Q=0$. Thus the main contribution to $\Sigma(K)$ comes from the vicinity of zero momentum for pairs, leading to the approximation of $\Sigma(K)\approx-\Delta^{2}G_{0}(-K)$, where we define the pseudogap as $\Delta^{2}=-\sum_{Q}t(Q)$. With this approximation, the full Green’s function takes a simple BCS-like form and is given by
@@
where $E_{\mathbf{k}}=\sqrt{\xi_{\mathbf{k}}^{2}+\Delta^{2}}$, $u_{\mathbf{k}}^{2}=(1+\xi_{\mathbf{k}}/E_{\mathbf{k}})/2$, and $v_{\mathbf{k}}^{2}=(1-\xi_{\mathbf{k}}/E_{\mathbf{k}})/2$. Then one obtains a generalized BCS-like gap equation,
@@
To extract the pair dispersion, one can Taylor expand $t^{-1}$ near $Q=0$ and analytically continue to the real frequency axis, with $(\mathrm{i}\Omega_{l}\rightarrow\Omega+\mathrm{i}0^{+})$, so that
@@
$a_{0}\Delta^{2}=\sum_{\mathbf{q}}\left[1+4\frac{a_{1}}{a_{0}}(\Omega_{\mathbf{q}}^{0}-\mu_{\text{p}})\right]^{-1/2}b(\Omega_{\mathbf{q}}),$ (4)
where $b(x)$ is the Bose distribution function and $\Omega_{\mathbf{q}}=\left[\sqrt{a_{0}^{2}+4a_{0}a_{1}(\Omega_{\mathbf{q}}^{0}-\mu_{\text{p}})}-a_{0}\right]/2a_{1}$ represents the pair dispersion. The coefficients $a_{0}$, $a_{1}$, $M_{\text{B}}$ are determined via the expansion process. Except in the weak coupling BCS regime, the $a_{1}$ term serves as a small quantitative correction and can often be neglected, so that $\Omega_{\mathbf{q}}\approx\Omega_{\mathbf{q}}^{0}-\mu_{\text{p}}$.

### v2

@@ Abstract
Proper treatment of the many-body interactions is of paramount importance in our understanding of strongly correlated systems. Here we investigate the effects of particle-hole fluctuations on the Berezinskii-Kosterlitz-Thouless (BKT) transition in two-dimensional Fermi gases throughout the entire BCS-BEC crossover. We include self-consistently in the self energy treatment the entire particle-hole $T$ matrix, which constitutes a renormalization of the bare interaction that appears in the particle-particle scattering $T$ matrix, leading to a screening of the pairing interaction and hence a reduction of the pairing gap and the transition temperature $T_{\text{BKT}}$. Here $T_{\text{BKT}}$ is determined by the critical phase space density, for which the pair density and pair mass are determined using a pairing fluctuation theory, which accommodates self-consistently the important self-energy feedback in the treatment of finite-momentum pairing fluctuations. The screening strength varies continuously from its maximum in the BCS limit to essentially zero in BEC limit. In the unitary regime, it leads to an interaction-dependent shift of $T_{\text{BKT}}$ towards the BEC regime. This shift is crucial in explaining experimental data, when quantitativeness is important, as they often depend on the interaction strength. Where experimental and simulation data are available, our findings are consistent with experiment in the unitary and BEC regimes and with quantum Monte Carlo simulations in the BCS and unitary regimes.
@@ I Introduction
Strongly correlated systems constitute a main challenge and are thus at the heart and frontier of condensed matter physics. Due to the low dimensionality, fluctuations in two dimensions (2D) are usually strong, positioning the system in the strongly correlated regime, for which a proper treatment of interaction effects is of paramount importance, in order to understand various experiments and physical phenomena at a quantitative level.
@@
In the absence of long range order à la the Mermin-Wagner theorem [36], phase transition in 2D is in general of the Berezinskii [3], Kosterlitz and Thouless (BKT) [31] type. Indeed, superfluid transitions in ultracold atomic Bose gases have been experimentally found to exhibit a strong BKT nature in 2D [22, 18, 48]. BKT transitions have been of wide interest since the most important high $T_{c}$ superconductors are quasi-2D layered materials [32], for which the low dimensionality is favorable for the formation of the wide-spread pseudogap phenomena, which are arguably attributable to strong fluctuations [35, 16]. It is also the model commonly used for describing the superconducting transition in thin films [25, 52]. Recent exciting discoveries of (quasi-)2D superconductors [6, 1, 26, 7] have added to the interest in 2D phase transitions. The BKT nature of the superfluid transition in 2D atomic Fermi gases has also been experimentally confirmed [44, 38]. There are also, albeit not many, theoretical studies of the BKT superfluid transition in a 2D Fermi system for both weak and strong pairing interactions, e.g., Ref. [49, 46].
@@
In this paper, we report our study on the important yet often neglected effect of the particle-hole fluctuations on 2D Fermi gas superfluidity throughout the BCS-BEC crossover, to which similar studies has not been reported in the literature [46]. Similar to its 3D counterpart, we find that particle-hole fluctuations lead to a screening to the pairing interaction, causing a shift of the superfluid transition curve toward the BEC regime as a function of the pairing strength. This substantially suppresses $T_{c}$ and the gap in the BCS and unitary regimes, and has an important physical significance when comparing between theory and experiment.
@@
It is important to note that there has been a historical controversy surrounding 2D fermionic superfluids, dating back to Kosterlitz and Thouless [31], which primarily concerns observable signatures and the applicability of BKT physics. It is not straightforward to apply the BKT theory based on the XY model to fermionic superfluids. In the XY model, in which the superfluid density $n_{s}/m$ provides the phase stiffness, it is the phase fluctuations of the vortex type that dominate the destruction of the quasi-long range superfluid order and drive the system into a disordered normal state. Unlike the superfluids of true bosons, however, the superfluid density is also suppressed by the pair-breaking Bogoliubov quasiparticle excitations as the temperature $T$ increases. In addition, both the density and the mass of the bosons (i.e., fermion pair) now depend on the temperature and the pairing interaction strength, except in the deep BEC regime. Therefore, computing the superfluid density using a mean-field approximation would yield a constant value $n_{s}/m=n/m$ at $T=0$ (or $n/4m$ from the boson point of view, with $n_{\text{B}}=n/2$ and $M_{\text{B}}=2m$), independent of the interaction strength. (Following standard notations, here $n$, $m$, $n_{\text{B}}$, $M_{\text{B}}$ and $n_{s}$ denote total fermion number density, fermion mass, Cooper pair number density, Cooper pair mass, and superfluid number density, respectively.) At finite $T$, the pair-breaking effect due to quasiparticle excitations can be partly accounted for via solving the mean-field BCS gap equation, leading to a mean-field result of $(n_{s}/m)^{\text{BCS}}$. However, not being able to properly take care of the effective pair mass and pair number density is thus expected to give rise to an overestimate of the superfluid transition temperature $T_{\text{BKT}}$. Indeed, the fact $T_{\text{BKT}}$, as measured experimentally in 2D atomic Fermi gases, varies with interaction strength, reflects that both $n_{\text{B}}$ and $M_{\text{B}}$ vary with the interaction.
@@
The pair number density at given temperature and interaction is normally governed by the pair dispersion. The latter, or equivalently pair mass, can, in principle, be extracted from the pair propagator. However, there are not many calculations of the fermionic $T_{\text{BKT}}$ in the literature [5, 2, 4, 37], partly due to the inapplicability of the simple XY model directly to the Fermi system. The fact that $(n_{s}/m)^{\text{BCS}}$ decreases with temperature below $T_{\text{BKT}}$ in the BCS regime manifests that amplitude fluctuations play an important role. Therefore, the superfluid transition in a 2D Fermi gas has a much richer physics than the simple XY model can capture. To account for the finite temperature and interaction effects, Wu et al [50] and Wang et al Ref. [49] determined $T_{\text{BKT}}$ using the critical phase space density criterion, based on a quantum Monte Carlo simulation [41] along with experimental support [38]. A pairing fluctuation theory [12] that was developed to address the pseudogap phenomena in 3D superconductors was used to determine the effective pair number density $n_{\text{B}}$ and pair mass $M_{\text{B}}$, which reflect the effects of both amplitude and phase fluctuations, governed by temperature and interaction strength. Nevertheless, these earlier works [49, 12] considered only the particle-particle channel of the $T$-matrix.
@@
In 2D, there have been no similar studies in the context of BCS-BEC crossover in the literature, however, except for the lowest order particle-hole contributions in the BCS limit [39]. It is the purpose of the present work to investigate the effect of particle-hole fluctuations on the Fermi gas superfluidity in 2D. The low dimensionality further enhances the strong pairing fluctuations, which are already present when the interaction is strong. The very BKT nature of the transition requires that there is already a sizable pairing amplitude, i.e., (pseudo)gap, at $T_{\text{BKT}}$. Therefore, this necessitates the self-consistent inclusion of the self-energy feedback in (both the particle-particle and) the particle-hole $T$-matrices. Here we incorporate the particle-hole channel contributions in the pairing fluctuation theory [12, 13, 17], and thus go beyond previous studies [50, 49]. This pairing fluctuation theory includes self-consistently the finite-momentum pairing fluctuations in the single-fermion self energy, and has successfully addressed multiple high $T_{c}$ [12, 13, 8, 9] and atomic Fermi gas experiments [15, 29, 14, 10]. In particular, it features naturally a pseudogap in 3D unitary Fermi gases, which has been unequivocally corroborated by a recent experiment [34]. Following the previous work in 3D [17], we show that the particle-hole $T$-matrix serves as a renormalized pairing interaction, which is to appear in the $T$-matrix in the usual particle-particle channel. We find that the particle-hole fluctuations lead to a screening to the pairing interaction, and the inverse pairing interaction is effectively shifted by the temperature-dependent, angular-averaged (at two different levels) particle-hole susceptibility $\langle\chi_{\text{ph}}\rangle$, which approaches a negative constant, $m/2\pi$, in the BCS limit and increases gradually in the crossover regime toward zero in the BEC limit. In result, the particle-hole channel shifts the $T_{\text{BKT}}$ curve towards the BEC regime as a function of pairing strength. Furthermore, comparison shows that the inclusion of particle-hole channel leads to a better overall agreement between our calculated $T_{\text{BKT}}$ with the results from experiment and quantum Monte Carlo simulations.
@@ II.1 Overview of the Pair Fluctuation Theory without the Particle-Hole Channel
We consider a 2D Fermi gas with a short-range $s$-wave attractive interaction $V_{\mathbf{k},\mathbf{k}^{\prime}}=U<0$, as described by a generic grand canonical Hamiltonian [13], which includes pairing with a finite center-of-mass momentum $\mathbf{q}$. The free fermion dispersion is given by $\xi_{\mathbf{k}}=\epsilon_{\mathbf{k}}-\mu\equiv\mathbf{k}^{2}/2m-\mu$, with chemical potential $\mu$. The Fermi momentum is $\hbar k_{\text{F}}=\sqrt{2\pi n}\hbar$, and Fermi energy $E_{\text{F}}\equiv k_{\text{B}}T_{\text{F}}=\hbar^{2}k_{\text{F}}^{2}/2m$. As usual, we shall use the natural units, and set $\hbar=k_{\text{B}}=1$ and the volume to unity [12].
The fermions acquire self energy via pair binding and unbinding. Our approximated equations, derived via an equations of motion approach [27, 11], is often cast into a $T$ matrix formalism, with a mix of bare Green’s function $G_{0}(K)=(\mathrm{i}\omega_{l}-\xi_{\mathbf{k}})^{-1}$ and full Green’s function $G(K)$ in the pair susceptibility $\chi(Q)=\sum_{K}G(K)G_{0}(Q-K)$. Here we use the four-momentum notation, $K\equiv(\mathrm{i}\omega_{n},\mathbf{k})$ and $Q\equiv(\mathrm{i}\Omega_{l},\mathbf{q})$, $\sum_{K}\equiv T\sum_{n,\mathbf{k}}$ and $\sum_{Q}\equiv T\sum_{l,\mathbf{q}}$, with $\omega_{n}$ and $\Omega_{l}$ being Matsubara frequencies for fermions and bosons, respectively [19]. The $T$-matrix $t(Q)$ is given by $t(Q)=U/[1+U\chi(Q)]$, with the self energy
@@
The Thouless criterion (for superfluid transition in 3D) requires $t^{-1}(Q=0)=U^{-1}+\chi(0)=0$. In order to accommodate the absence of long range order in 2D, we generalize the Thouless criterion to allow for a very small but finite pair chemical potential $\mu_{\text{p}}$, $U^{-1}+\chi(0)=t^{-1}(0)\propto\mu_{\text{p}}$. In this way, $t(Q)$ is highly peaked around $Q=0$. Thus the main contribution to $\Sigma(K)$ comes from the vicinity of $Q=0$, leading to a pseudogap approximation, $\Sigma(K)\approx-\Delta^{2}G_{0}(-K)$, where we define the pseudogap as $\Delta^{2}=-\sum_{Q}t(Q)$. With this approximation, the full Green’s function takes a simple BCS-like form and is given by
@@
where $E_{\mathbf{k}}=\sqrt{\xi_{\mathbf{k}}^{2}+\Delta^{2}}$, $u_{\mathbf{k}}^{2}=(1+\xi_{\mathbf{k}}/E_{\mathbf{k}})/2$, and $v_{\mathbf{k}}^{2}=(1-\xi_{\mathbf{k}}/E_{\mathbf{k}})/2$. This leads to a BCS-like gap equation,
@@
To extract the pair dispersion, one Taylor-expands $t^{-1}$ near $Q=0$ and analytically continues to real frequency, with $(\mathrm{i}\Omega_{l}\rightarrow\Omega+\mathrm{i}0^{+})$, so that
@@
$n_{\text{B}}\equiv a_{0}\Delta^{2}=\sum_{\mathbf{q}}\left[1+4\frac{a_{1}}{a_{0}}(\Omega_{\mathbf{q}}^{0}-\mu_{\text{p}})\right]^{-1/2}b(\Omega_{\mathbf{q}}),$ (4)
where $b(x)$ is the Bose distribution function and $\Omega_{\mathbf{q}}=\left[\sqrt{a_{0}^{2}+4a_{0}a_{1}(\Omega_{\mathbf{q}}^{0}-\mu_{\text{p}})}-a_{0}\right]/2a_{1}$ represents the pair dispersion. The coefficients $a_{0}$, $a_{1}$, $M_{\text{B}}$ are determined via the expansion process. Except in the weak coupling BCS regime, the $a_{1}$ term serves as a small quantitative correction and can often be neglected, so that $\Omega_{\mathbf{q}}\approx\Omega_{\mathbf{q}}^{0}-\mu_{\text{p}}$. The way Eq. (4) yields $n_{\text{B}}$ can be seen from the relation $n_{\text{B}}=n/2-\sum_{\mathbf{k}}f(\xi_{\mathbf{k}})$ [11].